a+5=b
a+5-5=b-5 is this a correct equation or incorrect equation and tell why

Answer by reynard2007(52)


This is correct because you're basically decreasing both sides of the equation by the same value. And both sides of the equation should be equal. Think that if x = y, then x-5 = y-5. If 9 = 9, then 9 - 5 = 9 - 5.